Part of $11,000 was invested by an investment club at 10% interest and the rest at 12%. The annual income from the investments is $1,260. How much was invested at each rate?
Answer by TimothyLamb(4379)

x = amount at 10%
y = amount at 12%
---
x + y = 11000
0.10x + 0.12y = 1260
---
put the system of linear equations into standard form
---
x - y = 49
x - 4y = -8
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = amount at 10% = 3000
y = amount at 12% = 8000
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---